OpenStudy (anonymous):

If h(x)=(fog) (x) and h(x)=3(x+2)^2, find f(x) and g(x)

OpenStudy (aum):

h(x)=3(x+2)^2 We need to split it into f(g(x)) In 3(x+2)^2 we see a squaring function and we see an adding function. so g(x) can be (x+2) and f(x) can be 3x^2

geerky42 (geerky42):

\((f\circ g)(x) = f(g(x))\) We can easily let g(x) = x, so f(x) can be equal to h(x) Since \(h(x) = 3(x+2)^2\), we can let \(f(x) = 3(x+2)^2\) So \(f(g(x)) = f(x) = 3(x+2)^2\) So we have \(f(x) = 3(x+2)^2\) and \(g(x) = x\) LOL at this question... very bypass-able haha

geerky42 (geerky42):

seriously, do what @aum did.

OpenStudy (anonymous):

Thanks you so much!

OpenStudy (aum):

You are welcome.